OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

So where do you think you should start?

OpenStudy (anonymous):

we will multiply 2 and 4 in parenthisis first : 2x+12-10= 12+4x-4 put the variables togathar and integers togather and add or substract them: 2x+2-8=0 2x+6=0 2x=-6 x=-3

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

2x+12-10 = 12+4x-4 2-12+4 = 4x-2x -6 = 2x x = -6/2 x = -3